OpenStudy (tomboy213):

please help! I have the answer, I just need to explain it. The table shows the height of a plant as it grows. What equation in point-slope form gives the plant’s height at any time? Let y stand for the height of the plant in cm and let x stand for the time in months. Time (weeks): 2; Plant Height (inches): 12 Time (weeks): 4; Plant Height (inches): 17 Time (weeks): 6; Plant Height (inches): 22 Time (weeks): 8; Plant Height (inches): 27

OpenStudy (tomboy213):

The answer is y=(5/2)x+7

ganeshie8 (ganeshie8):

Hi

OpenStudy (tomboy213):

Hi

563blackghost (563blackghost):

Heyo ^^ So you just need an explanation of how you found the answer?

OpenStudy (tomboy213):

yes

563blackghost (563blackghost):

Well if that is what you need to find we would convert these times into coordinates so.... \(\Huge{(2,12),~ (4,17), ~ (6,12),~(8,27)}\) So to find the slope we would choose two points and find the slope...So we would use the slope intercept formula... \(\Huge{\frac{y_{2}-y_{1}}{x_{2}-x_{1}}=\frac{17-12}{4-2}}\) In which you would simply simplify to find the slope in which what you got was \(\LARGE{\frac{5}{2}}\)....

OpenStudy (tomboy213):

Thank you blackghost, I apprectiate it!

563blackghost (563blackghost):

No Problem! Are you able to write an explanation to find the y-intercept?

OpenStudy (tomboy213):

Yes

563blackghost (563blackghost):

Great! Good luck with your work and have a great summer ^^

OpenStudy (tomboy213):

thank you, you too!